The computer-science aptitude score, x, and the achievement score, y (measured by a comprehensive final), were measured for 10 students in a beginning computer-science course. The results were as follows. Calculate the estimated standard error of regression, sb1. (Give your answer correct to two decimal places.)

x 4 16 20 13 22 21 15 20 19 16
y 20 18 23 35 26 25 26 29 18 26
Answer #1

SSE =Syy-(Sxy)2/Sxx= 242.041
s2 =SSE/(n-2)= 30.2552
std error σ              = =se =√s2= 5.5005
estimated std error of slope =se(β1) =s/√Sxx=0.3462 ~   0.35